Calculate the internal energy and enthalpy of 1 kg of air occupying 0.05m3 at 20 bar. If the internal energy is increased by 120 kJ/kg as air is compressed to 50 bar, calculate the new volume occupied by 1 kg of air. Assume air is a perfect gas



1
Expert's answer
2021-07-07T08:39:32-0400

Solution.

"m=1kg;"

"V_1=0.05m^3;"

"p_1=2\sdot10^6Pa;"

"\Delta U=120\sdot10^3J;"

"p_2=5\sdot10^6Pa;"

"M=28.96\sdot10^{-3}kg/mol;"

"T_1=293K;"

"U=\dfrac{m}{M}C_VT;"

"C_V=\dfrac{i}{2}R=\dfrac{5}{2}\sdot8.31=20.8J/molK;"

"U=\dfrac{1}{28.96\sdot10^{-3}}\sdot20.8\sdot293=210442J;"

"H=U+pV;"

"H=210442+2\sdot10^6\sdot0.05=310442J;"

"\Delta U=\dfrac{m}{M}C_V(T_2-T_1)\implies T_2=\dfrac{\Delta U M}{mC_V}+T_1;"

"T_2=\dfrac{120\sdot10^3\sdot28.96\sdot10^{-3}}{1\sdot20.8}+293=460K;"

"\dfrac{P_1\sdot V_1}{T_1}=\dfrac{P_2\sdot V_2}{T_2}\implies V_2=\dfrac{P_1V_1T_2}{T_1P_2};"

"V_2=\dfrac{2\sdot10^6\sdot0.05\sdot460}{293\sdot5\sdot10^6}=0.031m^3;"

Answer: "U=210442J;"

"H=310442J;"

"V_2=0.031m^3."